OpenStudy (anonymous):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 3/4 / 2 1/2 A. 8/2^3 B. /2^3/4 C.4/2 D. /2

Nnesha (nnesha):

\[\huge\rm \frac{ 2^\frac{ 3 }{ 4 } }{ 2^\frac{ 1 }{ 2 } }\] apply exponent rule \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] bases are same so you can move 1/2 exponent to the top when you do that sign would change

OpenStudy (anonymous):

C?

OpenStudy (jacob902):

Rewrite the rational exponent as a radical by extending the properties of integers exponents. 2^3/4 / 2^1/2 the choices are 8 radical2^3 Radical 2^3/4 4 radical 2 radical2

OpenStudy (anonymous):

ok thnx :)